OpenStudy (anonymous):

(-2,-1) & (-9,-4) slope intercept form

OpenStudy (anonymous):

Find slope: m = (y2 - y1) / (x2 - x1) Plug slope (m) into y = mx + b Plug (-2, -1) temporarily into x and y, leaving b as the only unknown. Solve for b (simple algebra). You now know m and b. Plug them into y = mx + b, using plain x and y for x and y (in other words, not using the -2 and -1 any more). Okay?

OpenStudy (anonymous):

Idon't know simple algebra because I have never had algebra. It has been 40 years since I was in school and all I had was basic math. This is very hard for me thus why I am here. I am trying thou.